# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ 6.2. $K_{LS}^{+}\leq K_{IS}^{+}$ in Inequality (k)
(6.4) $\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}\geq K_{LS}^{+}.$
Let $\alpha>0$. For each $a_{i}$, one can find an integer $n_{i}$ such that $e^{-(n_{i}+1)\alpha}<a_{i}\leq e^{-n_{i}\alpha}$. As $i\to\infty$, it holds that $n_{i}\to\infty$. So, by the monotonicity of $t\mapsto\frac{I_{n}(t)}{t}$ [32, Proposition 3.1] and $t\mapsto\sqrt{\ln\frac{1}{t}}$,
@@
$\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}$ $\geq\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2(n_{i}+1)\alpha}}$
$=\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}$
(6.5) $\geq\liminf_{i\to 0}\frac{I_{n_{i}}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}$
$\geq\sqrt{\frac{\breve{\Theta}(\alpha)}{2\alpha}},$
where (6.5) is due to that $I_{n}(a)\geq I_{m}(a)$ for all $m\geq n$ and any $a$. Since $\alpha>0$ is arbitrary, we obtain (6.4). ∎
6.3. Statement 3
We now prove Statement 3 in Theorem 2. For brevity, we denote $K_{1}:=K_{IS,1}^{+}=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}$. For any log-concave probability measure $\nu$ on $\mathbb{R}$, Bobkov [7] showed that $I(a)=f(F^{-1}(a))\wedge f(F^{-1}(1-a))$.
@@ 6.3. Statement 3
$\frac{f(F^{-1}(a))}{I_{G}(a)}$ $=\frac{f(F^{-1}(a))}{\varphi(\Phi^{-1}(a))}$
$\sim\frac{V^{\prime}(F^{-1}(a))}{\Phi^{-1}(a)}$
$\sim\frac{V^{\prime\prime}(F^{-1}(a))\varphi(\Phi^{-1}(a))}{f(F^{-1}(a))}$
$=V^{\prime\prime}(F^{-1}(a))\Big/\frac{I_{G}(a)}{I(a)},$
which implies
$\left(\frac{f(F^{-1}(a))}{I_{G}(a)}\right)^{2}$ $\sim V^{\prime\prime}(F^{-1}(a)).$
Similarly,
$\left(\frac{f(F^{-1}(1-a))}{I_{G}(a)}\right)^{2}$ $\sim V^{\prime\prime}(F^{-1}(1-a)).$
Hence,
$\left(\frac{I(a)}{I_{G}(a)}\right)^{2}$ $\sim V^{\prime\prime}(F^{-1}(a))\wedge V^{\prime\prime}(F^{-1}(1-a)),$
which implies
$K_{1}$ $=\lim_{x\to\infty}V^{\prime\prime}(x)\wedge\lim_{x\to-\infty}V^{\prime\prime}(x).$
@@
$V^{\prime}(x_{0})+(1-\epsilon)K_{1}(x-x_{0})$
$\leq$ $V^{\prime}(x)$
(6.6) $\leq$ $V^{\prime}(x_{0})+(1+\epsilon)K_{1}(x-x_{0}),$
@@
$\leq$ $V(x)$
(6.7) $\leq$ $V(x_{0})+V^{\prime}(x_{0})(x-x_{0})+\frac{(1+\epsilon)K_{1}}{2}(x-x_{0})^{2},$
@@
(6.8) $\int V(x+b)g(x){\text{\rm d}}x-\frac{1}{2}\ln(2\pi e/K_{1})\leq\frac{1}{2K}\int(V^{\prime}(x+b)-K_{1}x)^{2}g(x){\text{\rm d}}x.$
Substituting (6.6) and (6.7) into (6.8) and noting that both the parts $\int_{x_{0}}^{\infty}$ of the integrals at two sides can be negligible when $b$ is choosen to approach -$\infty$, we obtain
$\int\left(V(x_{0})+V^{\prime}(x_{0})(x+b-x_{0})+\frac{(1-\epsilon)K_{1}}{2}(x+b-x_{0})^{2}\right)g(x){\text{\rm d}}x-\frac{1}{2}\ln(2\pi e/K_{1})$
$\leq$ $\frac{1}{2K}\int\left(V^{\prime}(x_{0})+(1+\epsilon)K_{1}(x+b-x_{0})-K_{1}x\right)^{2}g(x){\text{\rm d}}x+o_{b}(1),$
@@
$V(x_{0})+V^{\prime}(x_{0})(b-x_{0})+\frac{(1-\epsilon)K_{1}}{2}\left(\frac{1}{K_{1}}+(b-x_{0})^{2}\right)-\frac{1}{2}\ln(2\pi e/K_{1})$
(6.9) $\leq$ $\frac{1}{2K}\left(\epsilon^{2}+\left(V^{\prime}(x_{0})+(1+\epsilon)K_{1}(b-x_{0})\right)^{2}\right)+o_{b}(1),$
where $o_{b}(1)$ is a term vanishing as $b\to$-$\infty$. Letting $b\to$-$\infty$ and comparing the coefficients of $b^{2}$ at two sides, we obtain
$\frac{(1-\epsilon)K}{(1+\epsilon)^{2}}\leq K_{1}.$
Noting that when $b\to$-$\infty$, the relative entropy (i.e., the LHS of (6.9)) approaches $\infty$. So, the constant $K$ can be set to $K_{LS}^{+}$. Letting $\epsilon\to 0$ yields the desired result.
7. Proof of Theorem 4
We next compare the dimension-free isoperimetric profile $I_{\inf}$ and the infinite-dimensional isoperimetric profile $I_{\infty}$. Recall that the dimension-free isoperimetric profile
$I_{\inf}(a):=\lim_{n\to\infty}I_{n}(a)=\inf_{n\geq 1}I_{n}(a),$
and $I_{\infty}$ is the isoperimetric profile for the space $(\mathcal{M}^{\infty},\nu^{\otimes\infty})$.
@@ 7. Proof of Theorem 4
We restrict the set $A$ in (7.1) to be closed, in order to ensure that $A^{r}\to A$ and hence $\nu_{n}(A^{r})\to\nu_{n}(A)$ as $r\downarrow 0$. This restriction usually does not affect the isoperimetric profile, e.g., for the Riemannian manifold setting with an absolutely continuous reference measure.
In the setting considered in this paper, the isoperimetric profiles $I_{n},n\in\mathbb{N}$ are all concave, which implies their pointwise minimum $I_{\inf}$ is concave as well and hence also continuous on $(0,1)$. So, Theorem 19 implies $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in(0,1)$. Moreover, obviously $I_{\inf}(a)=I_{\infty}(a)=0$ for $a\in\{0,1\}$.
Corollary 3.
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol},V\in C^{2}(\mathcal{M})$. Assume that (1.5) holds. Then, $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in[0,1]$.
This corollary is a restatement of Theorem 4. Hence it remains to prove Theorem 19.
7.1. Proof of Theorem 19
@@
Observe that $I_{\infty}(a)\leq I_{\inf}(a)$, since for any $n$ and $B\subseteq\mathcal{X}^{n}$, $B\times\mathcal{X}^{\mathbb{N}}$ in $\mathcal{X}^{\mathbb{N}}$ has the same measure and the same boundary as those of $B$ in $\mathcal{X}^{n}$. We next prove that $I_{\infty}(a)\geq I_{\inf}(a)$. We follow similar proof steps for Gaussian measures in [20, Theorem 2.2.4]. We first make the following claim.
@@
(b) Inclusion $K\subseteq\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))$: By definition, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$ for any $\mathbf{x}\in K$, so $\mathbf{x}\in\pi_{n}^{-1}(\pi_{n}(K))$ for all $n$. Thus,
$\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
Inclusion $\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))\subseteq K$: Suppose $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K))$. Then for each $n$, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$, so there exists $\mathbf{x}^{(n)}\in K$ such that $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$. Since $K$ is compact in the product topology of $\mathcal{X}^{\mathbb{N}}$, it is closed and sequentially compact. The sequence $(\mathbf{x}^{(n)})$ lies in $K$. The projections $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$ imply that the first $n$ coordinates of $\mathbf{x}^{(n)}$ match those of $\mathbf{x}$. This implies $\mathbf{x}^{(n)}\to\mathbf{x}$ in the product topology. Since $K$ is closed, $\mathbf{x}\in K$. Hence,
$K=\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
@@
which is compact in $X_{i}$ since $\pi_{i}$ is continuous. For any $\mathbf{y}\in K^{r}$, there exists $\mathbf{x}\in K$ such that
$\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}\leq r^{2}.$
Hence, for each coordinate $i$,
$d(y_{i},C_{i})\leq d(y_{i},x_{i})\leq r,$
because $x_{i}\in C_{i}$. This implies
@@
which is compact since $\mathcal{X}$ is locally compact. Therefore,
$K^{r}\subseteq\prod_{i=1}^{\infty}C_{i}^{r},$
and by Tychonoff’s theorem, $\prod_{i=1}^{\infty}C_{i}^{r}$ is compact in the product topology.
@@
$B_{r]}^{(N)}(\mathbf{x}):=\left\{\mathbf{y}:\sum_{i=1}^{N}d(x_{i},y_{i})^{2}\leq r^{2}\right\},$
and
$K_{N}^{r}=\bigcup_{\mathbf{x}\in K}B_{r]}^{(N)}(\mathbf{x}).$
@@
$\leq(\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},y_{i})^{2}}+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}})^{2}$
$\leq(r+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}})^{2},$
@@
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $=\lim_{n\to\infty}\frac{\nu_{\infty}(A_{n}^{r})-\nu_{\infty}(A_{n})}{r}$
$=\lim_{n\to\infty}\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r},$
@@
Since $f(t):=\nu_{n}(E_{n}^{r})$ is monotone, it is differentiable almost everywhere on $(0,r)$. Hence, by the mean value theorem, there is some $t_{n}\in(0,r)$ such that $f$ is differentiable at $t_{n}$, and moreover,
$\frac{f(r)-f(0)}{r}\geq f^{\prime}(t_{n}),$
i.e.,
$\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r}\geq\nu_{n}^{+}(E_{n}^{t_{n}}).$
Hence,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $\geq\limsup_{n\to\infty}\nu_{n}^{+}(E_{n}^{t_{n}})$
$\geq\limsup_{n\to\infty}I_{n}(\nu_{n}(E_{n}^{t_{n}}))$
$\geq\limsup_{n\to\infty}I_{\inf}(\nu_{n}(E_{n}^{t_{n}})).$
Observe that
$\nu_{\infty}(A_{n})=\nu_{n}(E_{n})\leq\nu_{n}(E_{n}^{t_{n}})\leq\nu_{n}(E_{n}^{r})=\nu_{\infty}(B_{n}).$
Taking $n\to\infty$,
$\nu_{\infty}(A)\leq\limsup_{n\to\infty}\nu_{n}(E_{n}^{t_{n}})\leq\nu_{\infty}(A^{r}).$
Therefore,
@@
Taking $r\downarrow 0$ and by the continuity of $I_{\inf}$ at $a$,
$\nu_{\infty}^{+}(A)$ $\geq I_{\inf}(a).$
So, Theorem 19 holds for compact sets.
@@
$\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A_{\epsilon})\leq a\leq\nu_{\infty}(A_{\epsilon}^{r})}I_{\inf}(a)$
$\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A)-\epsilon\leq a\leq\nu_{\infty}(A^{r})}I_{\inf}(a).$
@@
8. Proof of Proposition 2

### v3

@@ Appendix A-2. Inequality (l)
(6.15) $\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}\geq\sqrt{K_{LS}^{+}}.$
Let $\alpha>0$. For each $a_{i}$, one can find an integer $n_{i}$ such that $e^{-(n_{i}+1)\alpha}<a_{i}\leq e^{-n_{i}\alpha}$. As $i\to\infty$, it holds that $n_{i}\to\infty$. So, by the monotonicity of $t\mapsto\frac{I_{n}(t)}{t}$ [45, Proposition 3.1] and $t\mapsto\sqrt{\ln\frac{1}{t}}$,
@@
$\liminf_{i\to 0}\frac{I_{n}(a_{i})}{a_{i}\sqrt{2\ln\frac{1}{a_{i}}}}$ $\geq\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2(n_{i}+1)\alpha}}=\liminf_{i\to 0}\frac{I_{n}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}$
$\geq\liminf_{i\to 0}\frac{I_{n_{i}}(e^{-n_{i}\alpha})}{e^{-n_{i}\alpha}\sqrt{2n_{i}\alpha}}\geq\sqrt{\frac{\breve{\Theta}(\alpha)}{2\alpha}},$
where the second inequality is due to that $I_{n}(a)\geq I_{m}(a)$ for all $m\geq n$ and any $a$. Since $\alpha>0$ is arbitrary, we obtain (6.15).
Appendix A-3. Statement 2
We now prove Statement 2 in Theorem 2. For brevity, we denote $K_{1}:=K_{IS,1}^{+}=\lim_{a\to 0}\frac{I_{1}(a)}{a\sqrt{2\ln\frac{1}{a}}}$. For any log-concave probability measure $\nu$ on $\mathbb{R}$, Bobkov [11] showed that $I(a)=f(F^{-1}(a))\wedge f(F^{-1}(1-a))$.
@@ Appendix A-3. Statement 2
$\frac{f(F^{-1}(a))}{I_{G}(a)}$ $=\frac{f(F^{-1}(a))}{\phi(\Phi^{-1}(a))}\sim\frac{V^{\prime}(F^{-1}(a))}{\Phi^{-1}(a)}\sim\frac{V^{\prime\prime}(F^{-1}(a))\phi(\Phi^{-1}(a))}{f(F^{-1}(a))}$
which implies $\left(\frac{f(F^{-1}(a))}{I_{G}(a)}\right)^{2}\sim V^{\prime\prime}(F^{-1}(a)).$ Similarly, $\left(\frac{f(F^{-1}(1-a))}{I_{G}(a)}\right)^{2}\sim V^{\prime\prime}(F^{-1}(1-a)).$ Hence, $\left(\frac{I(a)}{I_{G}(a)}\right)^{2}\sim V^{\prime\prime}(F^{-1}(a))\wedge V^{\prime\prime}(F^{-1}(1-a)),$ which implies $K_{1}=\lim_{x\to\infty}V^{\prime\prime}(x)\wedge\lim_{x\to-\infty}V^{\prime\prime}(x).$
@@
(6.16) $V^{\prime}(x_{0})+(1-\epsilon)K_{1}(x-x_{0})\leq V^{\prime}(x)\leq V^{\prime}(x_{0})+(1+\epsilon)K_{1}(x-x_{0}),$
@@
(6.17) $\leq$ $V(x)\leq V(x_{0})+V^{\prime}(x_{0})(x-x_{0})+\frac{(1+\epsilon)K_{1}}{2}(x-x_{0})^{2},$
@@
(6.18) $\int V(x+b)g(x){\text{\rm d}}x-\frac{1}{2}\ln\frac{2\pi e}{K_{1}}\leq\frac{1}{2K}\int(V^{\prime}(x+b)-K_{1}x)^{2}g(x){\text{\rm d}}x.$
Substituting (6.16) and (6.17) into (6.18) and noting that both the parts $\int_{x_{0}}^{\infty}$ of the integrals at two sides are negligible when $b$ is chosen to approach -$\infty$, we obtain
$\int\left(V(x_{0})+V^{\prime}(x_{0})(x+b-x_{0})+\frac{(1-\epsilon)K_{1}}{2}(x+b-x_{0})^{2}\right)g(x){\text{\rm d}}x-\frac{1}{2}\ln\frac{2\pi e}{K_{1}}$
$\qquad\leq\frac{1}{2K}\int\left(V^{\prime}(x_{0})+(1+\epsilon)K_{1}(x+b-x_{0})-K_{1}x\right)^{2}g(x){\text{\rm d}}x+o_{b}(1),$
@@
$V(x_{0})+V^{\prime}(x_{0})(b-x_{0})+\frac{(1-\epsilon)K_{1}}{2}\left(\frac{1}{K_{1}}+(b-x_{0})^{2}\right)-\frac{1}{2}\ln\frac{2\pi e}{K_{1}}$
(6.19) $\qquad\leq\frac{1}{2K}\left(\epsilon^{2}+\left(V^{\prime}(x_{0})+(1+\epsilon)K_{1}(b-x_{0})\right)^{2}\right)+o_{b}(1),$
where $o_{b}(1)$ is a term vanishing as $b\to$-$\infty$. Letting $b\to$-$\infty$ and comparing the coefficients of $b^{2}$ at two sides, we obtain $\frac{(1-\epsilon)K}{(1+\epsilon)^{2}}\leq K_{1}.$ Noting that when $b\to$-$\infty$, the relative entropy (i.e., the LHS of (6.19)) approaches $\infty$. So, the constant $K$ can be set to $K_{LS}^{+}$. Letting $\epsilon\to 0$ yields the desired result.
Appendix B: Proof of Theorem 4
We next compare the dimension-free isoperimetric profile $I_{\inf}$ and the infinite-dimensional isoperimetric profile $I_{\infty}$. Recall that the dimension-free isoperimetric profile $I_{\inf}(a):=\lim_{n\to\infty}I_{n}(a)=\inf_{n\geq 1}I_{n}(a),$ and $I_{\infty}$ is the isoperimetric profile for the space $(\mathcal{M}^{\infty},\nu^{\otimes\infty})$.
@@
We restrict the set $A$ in (6.20) to be closed, in order to ensure that $A^{r}\to A$ and hence $\nu_{n}(A^{r})\to\nu_{n}(A)$ as $r\downarrow 0$. This restriction usually does not affect the isoperimetric profile, e.g., for the smooth Riemannian manifold setting with an absolutely continuous reference measure [48].
In our Riemannian manifold setting, under Assumption 1, the isoperimetric profiles $I_{n},n\in\mathbb{N}$ are all concave, which implies their pointwise minimum $I_{\inf}$ is concave as well and hence also continuous on $(0,1)$. So, Theorem 18 implies $I_{\inf}(a)=I_{\infty}(a)$ for all $a\in(0,1)$. Moreover, obviously $I_{\inf}(a)=I_{\infty}(a)=0$ for $a\in\{0,1\}$. Thus, $I_{\inf}$ is continuous on $[0,1]$. Theorem 4 is a consequence of Theorem 18. Hence it remains to prove Theorem 18.
Appendix B-1. Proof of Theorem 18
@@
Observe that $I_{\infty}(a)\leq I_{\inf}(a)$, since for any $n$ and $B\subseteq\mathcal{X}^{n}$, $B\times\mathcal{X}^{\mathbb{N}}$ in $\mathcal{X}^{\mathbb{N}}$ has the same measure and the same boundary as those of $B$ in $\mathcal{X}^{n}$. We next prove that $I_{\infty}(a)\geq I_{\inf}(a)$. We follow similar proof steps for Gaussian measures in [28, Theorem 2.2.4]. We first make the following claim.
@@
(b) Inclusion $K\subseteq\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))$: By definition, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$ for any $\mathbf{x}\in K$, so $\mathbf{x}\in\pi_{n}^{-1}(\pi_{n}(K))$ for all $n$. Thus, $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
Inclusion $\bigcap_{n}\pi_{n}^{-1}(\pi_{n}(K))\subseteq K$: Suppose $\mathbf{x}\in\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K))$. Then for each $n$, $\pi_{n}(\mathbf{x})\in\pi_{n}(K)$, so there exists $\mathbf{x}^{(n)}\in K$ such that $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$. Since $K$ is compact in the product topology of $\mathcal{X}^{\mathbb{N}}$, it is closed and sequentially compact. The sequence $(\mathbf{x}^{(n)})$ lies in $K$. The projections $\pi_{n}(\mathbf{x}^{(n)})=\pi_{n}(\mathbf{x})$ imply that the first $n$ coordinates of $\mathbf{x}^{(n)}$ match those of $\mathbf{x}$. This implies $\mathbf{x}^{(n)}\to\mathbf{x}$ in the product topology. Since $K$ is closed, $\mathbf{x}\in K$. Hence, $K=\bigcap_{n=1}^{\infty}\pi_{n}^{-1}(\pi_{n}(K)).$
@@
which is compact in $X_{i}$ since $\pi_{i}$ is continuous. For any $\mathbf{y}\in K^{r}$, there exists $\mathbf{x}\in K$ such that $\sum_{i=1}^{\infty}d(x_{i},y_{i})^{2}\leq r^{2}.$ Hence, for each coordinate $i$, $d(y_{i},C_{i})\leq d(y_{i},x_{i})\leq r,$ because $x_{i}\in C_{i}$. This implies
@@
which is compact since $\mathcal{X}$ is locally compact. Therefore, $K^{r}\subseteq\prod_{i=1}^{\infty}C_{i}^{r},$ and by Tychonoff’s theorem, $\prod_{i=1}^{\infty}C_{i}^{r}$ is compact in the product topology.
@@
$B_{r]}^{(N)}(\mathbf{x}):=\left\{\mathbf{y}:\sum_{i=1}^{N}d(x_{i},y_{i})^{2}\leq r^{2}\right\}\textrm{ and }K_{N}^{r}:=\bigcup_{\mathbf{x}\in K}B_{r]}^{(N)}(\mathbf{x}).$
@@
$\qquad\leq\Big(\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},y_{i})^{2}}+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}}\Big)^{2}\leq\Big(r+\sqrt{\sum_{i=1}^{m}d(x_{i}^{(N_{k})},z_{i})^{2}}\Big)^{2},$
@@
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $=\lim_{n\to\infty}\frac{\nu_{\infty}(A_{n}^{r})-\nu_{\infty}(A_{n})}{r}=\lim_{n\to\infty}\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r},$
@@
Since $f(t):=\nu_{n}(E_{n}^{r})$ is monotone, it is differentiable almost everywhere on $(0,r)$. Hence, by the mean value theorem, there is some $t_{n}\in(0,r)$ such that $f$ is differentiable at $t_{n}$, and moreover, $\frac{f(r)-f(0)}{r}\geq f^{\prime}(t_{n}),$ i.e., $\frac{\nu_{n}(E_{n}^{r})-\nu_{n}(E_{n})}{r}\geq\nu_{n}^{+}(E_{n}^{t_{n}}).$ Hence,
$\frac{\nu_{\infty}(A^{r})-\nu_{\infty}(A)}{r}$ $\geq\limsup_{n\to\infty}\nu_{n}^{+}(E_{n}^{t_{n}})\geq\limsup_{n\to\infty}I_{n}(\nu_{n}(E_{n}^{t_{n}}))\geq\limsup_{n\to\infty}I_{\inf}(\nu_{n}(E_{n}^{t_{n}})).$
Observe that $\nu_{\infty}(A_{n})=\nu_{n}(E_{n})\leq\nu_{n}(E_{n}^{t_{n}})\leq\nu_{n}(E_{n}^{r})=\nu_{\infty}(B_{n}).$ Taking $n\to\infty$, $\nu_{\infty}(A)\leq\limsup_{n\to\infty}\nu_{n}(E_{n}^{t_{n}})\leq\nu_{\infty}(A^{r}).$ Therefore,
@@
Taking $r\downarrow 0$ and by the continuity of $I_{\inf}$ at $a$, $\nu_{\infty}^{+}(A)\geq I_{\inf}(a).$ So, Theorem 18 holds for compact sets.
@@
$\qquad\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A_{\epsilon})\leq a\leq\nu_{\infty}(A_{\epsilon}^{r})}I_{\inf}(a)\geq-\frac{\epsilon}{r}+\inf_{\nu_{\infty}(A)-\epsilon\leq a\leq\nu_{\infty}(A^{r})}I_{\inf}(a).$
@@
Appendix C: Proof of Theorem 5
By an approximation argument of Otto and Villani [54, Appendix A], without affecting the optimal constant in (1.18), it suffices to consider $f\in W^{1,2}(\mathcal{M})$ satisfying the smoothness and boundedness conditions in (2.4). Without loss of generality, we can further assume $\mathbb{E}_{\nu}\left[f\right]=0$ and $\mathbb{E}_{\nu}\left[f^{2}\right]=1$, since otherwise we can do normalization on $f$ which neither affects the optimal constant in (1.18). Among all this kind of functions, in the following we let $f$ be the approximately optimal function attaining the optimal constant in (1.18), i.e., $\mathbb{E}[|\nabla f|^{2}]\leq(K_{P}+\delta)\cdot\mathrm{Var}_{\nu}(f)=K_{P}+\delta$ for $\delta>0$.
For such $f$, we consider the set $A_{t}=H_{t\sqrt{n}}(f)=\left\{\mathbf{x}:g_{n}(\mathbf{x})\leq t\right\}$, where $g_{n}(\mathbf{x})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}f(x_{i})$. We next determine the asymptotics of the volume and the perimeter of this set.
By the coarea formula,
(6.21) $\int_{|g_{n}-t|\leq\delta}|\nabla g_{n}|\,d\nu_{n}=\int_{t-\delta}^{t+\delta}\left(\int_{g_{n}=s}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\right){\text{\rm d}}s$
where $\varphi_{n}(\mathbf{x})=e^{-\sum_{i=1}^{n}V(x_{i})}$ is the density of $\nu_{n}$, ${\text{\rm d}}\mathrm{\mathrm{area}}$ is the induced volume form on the level set $\{g_{n}=s\}$, and $|\nabla g_{n}(\mathbf{x})|$ be the norm of the gradient of $g_{n}$ at the point $\mathbf{x}\in\mathcal{M}^{n}$. Thus, by the mean value theorem, the RHS above is larger than or equal to
(6.22) $2\delta\cdot\int_{g_{n}=t_{n}}\varphi_{n}(\mathbf{x})\,{\text{\rm d}}\mathrm{\mathrm{area}}(\mathbf{x})\geq 2\delta\cdot\nu_{n}^{+}(A_{t_{n}}),$
for some $t_{n}\in[t-\delta,t+\delta]$. Denoting $\mathbf{X}\sim\nu_{n}$, the LHS above is equal to
(6.23) $\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}\times\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right].$
By the central limit theorem, as $n\to\infty$,
(6.24) $\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}\to\Phi(t+\delta)-\Phi(t-\delta).$
We next determine the limit of $\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right]$. Note that $|\nabla g_{n}(\mathbf{X})|^{2}=\frac{1}{n}\sum_{i=1}^{n}|\nabla f|^{2}(X_{i})$. Let $\beta:=\mathbb{E}[|\nabla f|^{2}]\leq K_{P}+\delta.$ By the weak law of large number, $\lim_{n\to\infty}\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\}\to 1.$ By basic probabilistic inequalities, for $\delta>0$,
$\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta,|g_{n}(\mathbf{X})-t|\leq\delta\}$ $\geq\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}-\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|>\delta\},$
which implies
$\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\big||g_{n}(\mathbf{X})-t|\leq\delta\}$ $\geq 1-\frac{\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|>\delta\}}{\mathbb{P}\{|g_{n}(\mathbf{X})-t|\leq\delta\}}.$
By taking limit as $n\to\infty$, $\lim_{n\to\infty}\mathbb{P}\{\big||\nabla g_{n}(\mathbf{X})|^{2}-\beta\big|\leq\delta\big||g_{n}(\mathbf{X})-t|\leq\delta\}=1.$ Since $|\nabla f|^{2}$ is bounded by assumption (and thus, so is $|\nabla g_{n}|^{2}$), we obtain
(6.25) $\lim_{n\to\infty}\mathbb{E}\left[|\nabla g_{n}(\mathbf{X})|\mid|g_{n}(\mathbf{X})-t|\leq\delta\right]$ $\leq\sqrt{\beta+\delta}\leq\sqrt{K_{P}+\delta}+\delta.$
Combining (6.21), (6.22), (6.23), (6.24), and (6.25) yields
$\limsup_{n\to\infty}\nu_{n}^{+}(A_{t_{n}})\leq\frac{\Phi(t+\delta)-\Phi(t-\delta)}{2\delta}\left(\sqrt{K_{P}+\delta}+\delta\right).$
Letting $\delta\to 0$ yields
(6.26) $\limsup_{\delta\to 0}\limsup_{n\to\infty}\nu_{n}^{+}(A_{t_{n}})\leq\sqrt{K_{P}}\phi(t).$
By the central limit theorem, $g_{n}(\mathbf{X})$ converges in distribution to $Z\sim\mathcal{N}(0,1)$. Thus, the volume $\nu_{n}(A_{t_{n}})$ is asymptotically sandwiched between $\Phi(t-\delta)$ and $\Phi(t+\delta)$ as $n\to\infty$, which implies
(6.27) $\lim_{\delta\to 0}\lim_{n\to\infty}\nu_{n}(A_{t_{n}})=\Phi(t).$
Equations (6.26) and (6.27) implies that there is a sequence of sets $B_{n}\subseteq\mathcal{M}^{n}$ such that $\lim_{n\to\infty}\nu_{n}^{+}(B_{n})\leq\sqrt{K_{P}}\phi(t)$ and $\lim_{n\to\infty}\nu_{n}(B_{n})=\Phi(t)$. By continuity of $I_{\inf}$, it holds that $I_{\inf}(a)\leq\sqrt{K_{P}}\cdot I_{\mathrm{G}}(a)$.
Appendix D: Proof of Proposition 2